Joann is five years older than her sister Teresa, Joann's age is one less than twice Teresa's. Find the ages of both girls.
Answer by stanbon(26297)

Joann is five years older than her sister Teresa, Joann's age is one less than twice Teresa's. Find the ages of both girls.
------------------------
Equations:
j = t + 5
j = 2t - 1
------------------
Substitute and solve for "t":
t+5 = 2t-1
t = 6 (Teresa's age)
------------
Since j= t+5, j=11 (Joann's age)
================================
Cheers,
Stan H.